OpenStudy (anonymous):

Input the equation of the given line in standard form. The line with m = -1/2 and the point (-2, -3).

OpenStudy (saifoo.khan):

find the y-intercpt.

OpenStudy (anonymous):

y=-1/2x - 4

OpenStudy (anonymous):

point slope formula

OpenStudy (saifoo.khan):

y = mx + c -3 = -1/2 (-2) + c c = -4 thereofre: y = -1x/2 - 4